OpenStudy (anonymous):

6a = 5/7 A.5 / 13 B.30 / 7 C.5 / 42 D.11 / 7

OpenStudy (hoblos):

divide both sides by 6 a= (5/7)/6 = 5/42

OpenStudy (hoblos):

C

OpenStudy (anonymous):

b